# Principal component analysis of absorbing state phase transitions

### v1

@@ I Introduction
A recent and promising avenue to enhance our comprehension of phase transitions involves the use of data-driven techniques, especially in their application to experimental contexts [11, 12]. Notably, in recent years, both supervised and unsupervised machine learning techniques have demonstrated their capability to identify and characterize equilibrium phase transitions, even in systems featuring intricate non-local order parameters [13, 14, 15, 16, 17, 18, 19, 20]. Whether the efficacy of these techniques extends to systems that are out of equilibrium remains relatively unexplored, although some investigations in this direction have been recently pursued [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
@@ II Models
Lattice models play an important role in understanding equilibrium phase transition. The equilibrium properties of these models at inverse temperature $\beta$ are determined by an energy cost function, the Hamiltonian $H$, which characterizes the equilibrium distribution $P_{s}(\mathcal{C})\propto e^{-\beta H[\mathcal{C}]}$ of a configuration $\mathcal{C}$ of the system. Non-equilibrium lattice models, instead, are defined by some dynamical evolution rules or, equivalently, by some set of transition probabilities $\{W(\mathcal{C}\to\mathcal{C}^{\prime})\}$, which specify how the system evolves in time from configuration $\mathcal{C}$ to $\mathcal{C}^{\prime}$ [10]. Unlike their equilibrium counterparts, the stationary probability distribution $P_{s}(\mathcal{C})$ of the configurations of these non-equilibrium lattice models is not known a priori. Indeed, $P_{s}(\mathcal{C})$ can only be determined as the solution of a master equation, which in most of the cases of statistical systems cannot be solved exactly.
In our study, we consider two lattice models defined in one spatial dimension which are driven out of equilibrium by the presence of an absorbing state, i.e. a state which cannot be left by the dynamics. The two models feature the same configuration space and have the same microscopic absorbing state, but belong to different universality classes: directed bond percolation and branching and annihilating random walks with an even number of offspring. In both cases, the presence of an absorbing state implies that the system is out of equilibrium. Indeed, at equilibrium the transition rates from one configuration to another should satisfy the law of detailed balance $P_{s}(\mathcal{C})W(\mathcal{C}\to\mathcal{C}^{\prime})=P_{s}(\mathcal{C}^{\prime})W(\mathcal{C}^{\prime}\to\mathcal{C})$ for every pair of configurations $\mathcal{C}$ and $\mathcal{C}^{\prime}$. In the case of an absorbing state $\mathcal{C}_{\rm{abs}}$, the rate of leaving the absorbing configuration is $W(\mathcal{C}_{\rm{abs}}\to\mathcal{C}^{\prime})=0$ for every $\mathcal{C}^{\prime}$. For this reason, a system featuring an absorbing state cannot obey detailed balance, making it a genuinely out-of-equilibrium system. In the following, we specify the dynamics of the two models under consideration.
@@ II.0.2 Branching and Annihilating Random Walks
Another well-established nonequilibrium universality class is the parity conserving one. This universality class is the simplest possible deviation from the DP universality class and it emerges in various contexts, including branching and annihilating random walks [48], nonequilibrium Ising models [49], systems with two absorbing states [50, 51], roughening transitions [52], and entanglement dynamics [53, 54].
@@ II.0.3 Observables and Phase Transitions
$z$ $1.580745(10)$ $1.750(5)$
$\alpha$ $0.159464(6)$ $0.285(2)$
Table 1: Estimates of the values of the critical points $p_{c}$ for the directed bond percolation (DBP) and for the BARWe with $n=2$ [55] and of the universal critical exponents for the corresponding universality classes, i.e., respectively, directed percolation (DP) [8, 57, 58] and parity conserving (PC) (estimated from the $n=4$ BARW in Ref. [59]), in one spatial dimension. In $d=1$ both DP and PC have $\alpha=\delta$, where $\delta$ is the exponent characterizing the decay of the survival probability [9].
@@ III.0.1 Principal Component Analysis
$\widetilde{\mathbf{X}}_{\mu i}=\mathbf{X}_{\mu i}-\frac{1}{N_{r}}\sum^{N_{r}}_{\mu=1}\mathbf{X}_{\mu i}.$ (13)
@@ III.0.2 Datasets
A natural question that arises is whether the spectral properties of the covariance matrix $\mathbf{\Sigma}$ or of the second-moment matrix $\mathbf{M}$, which are naturally expected to be sensitive to such a change, can be used to characterize the kind of transition undergone by the system as $p$ is varied. In our examination below, we find that both kinds of analyses can be used to describe the presence of such a phase transition, with an explicit connection to physical observables in the uncentered case. In view of this connection, below we focus on the analysis of $\mathbf{M}$, while we present the results concerning $\mathbf{\Sigma}$ in Appendix C. Let us note, however, that it is possible to obtain connections to physical observables based on a centered PCA, as illustrated for DBP in Appendix D.
@@ V.1 Largest normalized eigenvalue
As we did for DBP, we first analyze the behavior of the largest normalized uncentered eigenvalue $\pi^{u}_{1}$ as a function of time, which is reported in Fig. 9 for three values of $p$ corresponding to the active, critical, and absorbing phase and with $L=512$. In this case, the analysis is complicated by the fact that in parity-conserving models the decay of the density $\rho(t)$ as a function of time $t$ is algebraic also within the absorbing phase [8]. Indeed, such algebraic decay is observed in Fig. 9 both for $\pi^{u}_{1}$ and $\rho_{S}$. The expected power-law decay in the absorbing phase is the one associated with the pair-annihilation fixed point in one dimension $\alpha=1/2$ (see Ref. [73, 74]) and it is shown in Fig. 9 as a dashed line. As a matter of fact, this power can be observed in numerical simulations only at very low values of $p$ or at very large system sizes and the decay exponent varies continuously for finite system sizes. For the sizes accessible to our numerical simulations (up to $L=512$), it turned out to be difficult to disentangle finite-size effects from a possible crossover in the algebraic exponent due to the uncertainties in the location of the critical value $p_{c}$ of the transition. As a consequence, we did not manage to extract a decay exponent for $\pi^{u}_{1}(t)$ which is compatible with the value expected for $\alpha$ within PC universality class. We notice that this is not a peculiarity of this case: such finite volume effects are typically observed in the presence of algebraic decay of correlations, and occur in equilibrium as well – one example being Berezinski-Kosterlitz-Thouless phase transition.
@@
To proceed further with our analysis, we also perform a finite-size scaling analysis of the time dependence of $\pi^{u}_{1}(t)$ during an evolution with $p=p_{c}$ (see Eq. (21)), reported in Fig. 12(b). The value $\omega_{*}=-0.459(3)$ extracted from this analysis is compatible with the value extracted above from the analysis of the steady-state data and thus does not agree with the value $-\beta/\nu_{\perp}$ expected for PC. Concerning $\phi_{*}$, instead, we find a value $\phi_{*}=1.74(3)$ compatible with the dynamical critical exponent $z=1.750(5)$ of PC universality class (see Tab. 1).
@@ VI Conclusions
In this case, the success of this approach can be attributed to the fact that $\pi^{u}_{1}$, even close to criticality, is well separated from the smaller eigenvalues. At the same time, we have shown that entropic quantities that keep track of the entire PCA spectrum can serve as indicators of criticality, as shown in the case of DBP (Fig. 7). For models showing a PCA spectrum with a richer structure, such as BARWe, the asimilarity between the scaling properties of the largest normalized eigenvalue and of the order parameter of the model is considerably weaker — a fact which we attribute to larger finite-volume effects. Nevertheless, some quantities, such as the critical point and the dynamical critical exponent, can still be inferred accurately from the analysis of the sole largest normalized eigenvalue $\pi^{u}_{1}$. Importantly, we elucidate our numerical observations by employing linear algebra arguments, which form the basis of PCA, thereby leveraging its interpretable nature.
@@ Appendix D Quantified principal components
The second QPC $\langle|\mathrm{PC}_{2}|\rangle$ in Fig. 23 (a) clearly signals the onset of the phase transition in a fashion that closely resembles the behavior of a susceptibility for a system at equilibrium featuring a continuous (second-order) phase transition. In order to test if this quantity has good scaling properties we perform a finite-size scaling with the ansatz

### v2

@@ I Introduction
A recent and promising avenue to enhance our comprehension of phase transitions involves the use of data-driven techniques, especially in their application to experimental contexts [11, 12]. Notably, in recent years, both supervised and unsupervised machine learning techniques have demonstrated their capability to identify and characterize equilibrium phase transitions, even in systems featuring intricate non-local order parameters [13, 14, 15, 16, 17, 18, 19, 20]. Whether the efficacy of these techniques extends to systems that are out of equilibrium remains relatively unexplored, although some investigations in this direction have been recently pursued [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
@@ II Models
Lattice models play an important role in understanding equilibrium phase transition. The equilibrium properties of these models at inverse temperature $\beta$ are determined by an energy function, the Hamiltonian $H$, which characterizes the equilibrium distribution $P_{s}(\mathcal{C})\propto e^{-\beta H[\mathcal{C}]}$ of a configuration $\mathcal{C}$ of the system. Non-equilibrium lattice models, instead, are defined by some dynamical evolution rules or, equivalently, by some set of transition probabilities $\{W(\mathcal{C}\to\mathcal{C}^{\prime})\}$, which specify how the system evolves in time from configuration $\mathcal{C}$ to $\mathcal{C}^{\prime}$ [10]. Unlike their equilibrium counterparts, the stationary probability distribution $P_{s}(\mathcal{C})$ of the configurations of these non-equilibrium lattice models is not known a priori. Indeed, $P_{s}(\mathcal{C})$ can only be determined as the solution of a master equation, which in most of the cases of statistical systems cannot be solved exactly.
In our study, we consider two lattice models defined in one spatial dimension which are driven out of equilibrium by the presence of an absorbing state, i.e., a state which cannot be left by the dynamics. The two models feature the same configuration space and have the same microscopic absorbing state, but belong to different universality classes: directed bond percolation and branching and annihilating random walks with an even number of offspring. In both cases, the presence of an absorbing state implies that the system is out of equilibrium. Indeed, at equilibrium the transition rates from one configuration to another should satisfy the condition of detailed balance $P_{s}(\mathcal{C})W(\mathcal{C}\to\mathcal{C}^{\prime})=P_{s}(\mathcal{C}^{\prime})W(\mathcal{C}^{\prime}\to\mathcal{C})$ for every pair of configurations $\mathcal{C}$ and $\mathcal{C}^{\prime}$. In the case of an absorbing state $\mathcal{C}_{\rm{abs}}$, the rate of leaving the absorbing configuration is $W(\mathcal{C}_{\rm{abs}}\to\mathcal{C}^{\prime})=0$ for every $\mathcal{C}^{\prime}$. For this reason, a system featuring an absorbing state cannot obey detailed balance [^1] We point out that the existence of an absorbing state $\mathcal{C}_{\textrm{abs}}$ causes a violation of the detailed balance conditions only if the stationary distribution $P_{s}(\mathcal{C})$ is not concentrated on $\mathcal{C}_{\textrm{abs}}$, i.e., if $P_{s}(\mathcal{C})\neq\delta_{\mathcal{C},\mathcal{C}_{\textrm{abs}}}$. This concentration actually occurs in systems of finite size and thus a genuine non-equilibrium state may emerge only in the thermodynamic limit. , making it a genuinely out-of-equilibrium system. In the following, we specify the dynamics of the two models under consideration.
@@ II.0.2 Branching and Annihilating Random Walks
Another well-established nonequilibrium universality class is the parity conserving one, which was first observed in Refs [49, 50]. This universality class is the simplest possible deviation from the DP universality class and it emerges in various contexts, including branching and annihilating random walks [51], nonequilibrium Ising models [52], systems with two absorbing states [53, 54], roughening transitions [55], and entanglement dynamics [56, 57].
@@ II.0.3 Observables and Phase Transitions
$z$ $1.580745(10)$ $1.7415(5)$
$\alpha$ $0.159464(6)$ $0.2872(1)$
Table 1: Estimates of the values of the critical points $p_{c}$ for the directed bond percolation (DBP) and for the BARWe with $n=2$ [58] and of the universal critical exponents for the corresponding universality classes, i.e., respectively, directed percolation (DP) [8, 60, 61] and parity conserving (PC) (the estimates for $\beta,\nu_{\bot},\nu_{\parallel}$ were obtained from the $n=4$ BARW in Ref. [62], while the estimates for $z$ and $\alpha$ are taken from [63] for the $n=2$ BARW), in one spatial dimension. In $d=1$ both DP and PC have $\alpha=\delta$, where $\delta$ is the exponent characterizing the decay of the survival probability [9].
@@ III.0.1 Principal Component Analysis
$\widetilde{\mathbf{X}}_{\mu i}=\mathbf{X}_{\mu i}-\frac{1}{N_{r}}\sum^{N_{r}}_{\mu^{\prime}=1}\mathbf{X}_{\mu^{\prime}i}.$ (13)
@@ III.0.2 Datasets
A natural question that arises is whether the spectral properties of the covariance matrix $\mathbf{\Sigma}$ or of the second-moment matrix $\mathbf{M}$, which are naturally expected to be sensitive to such a change, can be used to characterize the kind of transition undergone by the system as $p$ is varied. In our examination below, we find that both kinds of analyses can be used to describe the presence of such a phase transition, with an explicit connection to physical observables in the uncentered case. In view of this connection, below we focus on the analysis of $\mathbf{M}$, while we present the results concerning $\mathbf{\Sigma}$ in Appendix C. Let us note, in addition, that also a centered analysis based on the quantified principal components, defined and discussed in Appendix D for DBP, can be physically informative.
@@ V.1 Largest normalized eigenvalue
As we did for DBP, we first analyze the behavior of the largest normalized uncentered eigenvalue $\pi^{u}_{1}$ as a function of time, which is reported in Fig. 9 for three values of $p$ corresponding to the active, critical, and absorbing phase and with $L=512$. In this case, the analysis is complicated by the fact that in parity-conserving models the decay of the density $\rho(t)$ as a function of time $t$ is algebraic also within the absorbing phase [8]. Indeed, such algebraic decay is observed in Fig. 9 both for $\pi^{u}_{1}$ and $\rho_{S}$. The expected power-law decay in the absorbing phase is the one associated with the pair-annihilation fixed point in one dimension $\alpha=1/2$ (see Ref. [78, 79]) and it is shown in Fig. 9 as a dashed line. As a matter of fact, this power can be observed in numerical simulations only at very low values of $p$ or at very large system sizes and the decay exponent varies continuously for finite system sizes. For the sizes accessible to our numerical simulations (up to $L=512$), it turned out to be difficult to disentangle finite-size effects from a possible crossover in the algebraic exponent. Indeed, the system sizes used in our simulations are much smaller than those used, for example, in Ref. [63]. As a consequence, we did not manage to extract a decay exponent for $\pi^{u}_{1}(t)$ which is compatible with the value expected for $\alpha$ within PC universality class. We notice that this is not a peculiarity of this case: such finite volume effects are typically observed in the presence of algebraic decay of correlations, and occur in equilibrium as well – one example being Berezinski-Kosterlitz-Thouless phase transition.
@@
To proceed further with our analysis, we also perform a finite-size scaling analysis of the time dependence of $\pi^{u}_{1}(t)$ during an evolution with $p=p_{c}$ (see Eq. (21)), reported in Fig. 12(b). The value $\omega_{*}=-0.459(3)$ extracted from this analysis is compatible with the value extracted above from the analysis of the steady-state data and thus does not agree with the value $-\beta/\nu_{\perp}$ expected for PC. Concerning $\phi_{*}$, instead, we find a value $\phi_{*}=1.74(3)$ compatible with the dynamical critical exponent $1.7415(5)$ of PC universality class (see Tab. 1).
@@ VI Conclusions
In this case, the success of this approach can be attributed to the fact that $\pi^{u}_{1}$, even close to criticality, is well separated from the smaller eigenvalues. At the same time, we have shown that entropic quantities that keep track of the entire PCA spectrum can serve as indicators of criticality, as shown in the case of DBP (Fig. 7). For models showing a PCA spectrum with a richer structure, such as BARWe, the similarity between the scaling properties of the largest normalized eigenvalue and of the order parameter of the model is considerably weaker — a fact which we attribute to larger finite-volume effects. Nevertheless, some quantities, such as the critical point and the dynamical critical exponent, can still be inferred accurately from the analysis of the sole largest normalized eigenvalue $\pi^{u}_{1}$. Importantly, we elucidate our numerical observations by employing linear algebra arguments, which form the basis of PCA, thereby leveraging its interpretable nature.
@@ Appendix D Quantified principal components
The second QPC $\langle|\mathrm{PC}_{2}|\rangle$ in Fig. 23(a) clearly signals the onset of the phase transition in a fashion that closely resembles the behavior of a susceptibility for a system at equilibrium featuring a continuous (second-order) phase transition. Note that $\langle|\mathrm{PC}_{2}|\rangle$ reminds of a staggered density (see $\vec{w}_{2}$ in Fig. 22) and, to our knowledge, it has not been previously examined for the DBP. Accordingly, the fact discussed below that it displays an interesting scaling behavior can be considered a genuine discovery of the PCA. In order to test if this quantity has good scaling properties we perform a finite-size scaling with the ansatz